Find the standard equation of the parabola which satisfies the given conditions.

1. Focus (−2, −5), directrix x = 6


2. Vertex (−4,2), focus (−4, −1)


3. Vertex (−8,3), directrix x = −10.5


4. Focus (7,11), directrix y = 4


1
Expert's answer
2021-03-18T04:01:11-0400

1. The distance of the directrix x=6 to the focus is 8 units therefore, 2p=-8, p=-4. Thus, the vertex is (2,-5).


The axis of symmetry is parallel to the x-axis:

(y - k)2 = 4p(x - h) h=2 and k=-5

(y-(-5))2= 4(-4)(x-2)


(y+5)2 = -16(x-2)

2. "(x-h)^2=4p(y-k)"

h=-4

k=2

p=-1-2=-3

"(x+4)^2=-12(y-2)"

"x^2+8x+16=-12y+24"

"y=-x^2\/12-2\/3x+2\/3"

3. (y - k)2= 4p(x - h) h=-8 and k=3

k-p=-10.5

p=10.5-8=2.5

"(y-3)^2=10(x+8)"

4."\sqrt{(y-k)^2 }=\sqrt{(x-a)^2+(y-b)^2}"

k=4

a=7

b=11

"\sqrt{(y-4)^2 }=\sqrt{(x-7)^2+(y-11)^2}"

"(y-4)^2 =(x-7)^2+(y-11)^2"

"y^2-8y+16=x^2-14x+49+y^2-22y+121"

"14y=x^2-14x+154"

"y=x^2\/14-x+11"